OpenStudy (anonymous):

Marty has a piggy bank that contains $2.25. The piggy bank contains only quarters and dimes. There are twice as many dimes as quarters in the bank.How many dimes are in the piggy bank?

OpenStudy (anonymous):

I needed to know how many dimes not quarters:(

OpenStudy (anonymous):

paige!!!!!!! i gave u the solution

OpenStudy (anonymous):

What that the dimes are two and the quarters are five?

OpenStudy (anonymous):

Please just help me?

Parth (parthkohli):

Dime is 10 cents

OpenStudy (anonymous):

Ohh so the answer is ten?

OpenStudy (anonymous):

or one dime?

Parth (parthkohli):

20x + 25x = 225 45x = 225 x = 225/45 x= 5 there were 2x dimes = 10 dimes and x quarters = 5 MARTY HAD 10 DIMES AND 5 QUARTERS in his piggy bank

OpenStudy (anonymous):

ok...

Parth (parthkohli):

So, he had 10*10 = 100 cents + 25 * 5 = 125 cents. That makes 2.25 dollars. We have proved the answer.